# Double-pair Coulomb(-Breit) photon correction to the correlated relativistic energy

### v1

@@ Abstract
The simplest, algebraic quantum-electrodynamical corrections, due to the double-negative energy subspace and instantaneous interactions, are computed to the no-pair Dirac-Coulomb(-Breit) energy of two-spin-1/2-fermion systems. Numerical results are reported for two-electron atoms with a clamped nucleus and positronium-like genuine two-particle systems. The Bethe-Salpeter equation provides the theoretical framework, and numerical methods have been developed for its equal-time time-slice. In practice, it requires solving a sixteen-component eigenvalue equation with a two-particle Dirac Hamiltonian, including the appropriate interaction. The double-pair corrections can either be included in the interaction part of the eigenvalue equation or treated as a perturbation to the no-pair Hamiltonian. The numerical results have an $\alpha$ fine-structure constant dependence that is in excellent agreement with the known $\alpha^{3}E_{\mathrm{h}}$-order double-pair correction of non-relativistic quantum electrodynamics.
@@ I Introduction
In a series of recent work Mátyus et al. (2023); Jeszenszki et al. (2021, 2022); Ferenc et al. (2022a); Ferenc and Mátyus (2023); Jeszenszki and Mátyus (2023); Margócsy and Mátyus (2024); Hollósy et al. (2024); Nonn et al. (2024), we have demonstrated that it should be possible, starting from the Bethe-Salpeter equation and its equal-time time-slice, to define a relativistic QED approach, in which a correlated relativistic reference state is first (numerically) computed to high precision Jeszenszki et al. (2021, 2022); Ferenc et al. (2022a); Ferenc and Mátyus (2023), and then, quantum electrodynamical corrections are computed to it by perturbation theory Mátyus et al. (2023); Margócsy and Mátyus (2024); Nonn et al. (2024). This second step is still in an early phase of research, and this work is one of the first steps to report actual methodologies and numerical values for the simplest QED corrections to a correlated relativistic reference state. This work considers the simplest, algebraic (non-energy-dependent) correction. When further, important (but more complicated) corrections, i.e., due to retardation, multi-photon, and radiative, self-energy, vertex, and vacuum polarization ‘effects’ will be computed Mátyus et al. (2023); Margócsy and Mátyus (2024); Nonn et al. (2024), then, this direction will complement the already existing $1/Z$ and nrQED approaches, and may help solve the helium puzzle or provide independent tests for (parts of) the other approaches.
@@ II Double-pair ladder correction to the no-pair Dirac-Coulomb(-Breit) energy
$\Phi(\bm{r}_{1},\bm{r}_{2})$ is the equal-time time-slice of the two-particle wave function, depending only on the particles’ spatial coordinates, $\bm{r}_{1}$ and $\bm{r}_{2}$. $H_{\text{np}}$ is the no-pair Dirac-Coulomb(-Breit) Hamiltonian and the $V_{\delta}+H_{\Delta}(E)$ term carries the corrections beyond the no-pair processes. The present work focuses on instantaneous interactions; hence, we drop the energy-dependent $H_{\Delta}(E)$ term and will work with the algebraic $V_{\delta}$ contribution.
@@
$V_{\delta}$ $=-L_{--}V_{\text{i}}L_{--}+L_{++}V_{\text{i}}L_{--}-L_{--}V_{\text{i}}L_{++}\ .$ (7)
$V_{\delta}$ couples the positive- (++) and negative- (--) energy subspaces through the $V_{\text{i}}$ instantaneous interaction.
@@
and then, $V_{\delta}$ is accounted for by perturbation theory.
For solutions of the no-pair wave equation, Eq. (8), it is convenient to explicitly write out the contributions of the $\Phi_{\text{np},i}$ according to the $\mathcal{L}_{++}$, $\mathcal{L}_{\text{BR}}=\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$, and $\mathcal{L}_{--}$ subspaces,
@@
in other words, $\phi_{\text{np},i}^{x}=L_{x}\Phi_{\text{np},i}L_{x}$ ($x=++,\text{BR}$ or --). The no-pair Hamiltonian is block diagonal over these subspaces $\mathcal{L}_{++},\mathcal{L}_{\text{BR}},$ and $\mathcal{L}_{--}$ subspaces. The non-trivial two-particle solutions, which correspond to the physically relevant bound and excited states, are in the $\mathcal{L}_{++}$ positive-energy subspace, hence we can focus on solving the eigenvalue equation on this subspace,
@@
where $E_{m}^{--}=E_{\text{np},i}$ with $i\in\mathcal{I}_{--}$ and $\mathcal{I}_{--}$ collects the indexes of the negative-energy solutions of Eq. (8).
@@
$E_{p}^{(1)}$ $=\langle\phi^{++}_{\text{np},p}|V_{\delta}|\phi^{++}_{\text{np},p}\rangle$ (15)
$=\langle\phi^{++}_{\text{np},p}|[-{L}_{--}V_{\text{i}}{L}_{--}+{L}_{++}V_{\text{i}}{L}_{--}-{L}_{--}V_{\text{i}}{L}_{++}]|\phi^{++}_{\text{np},p}\rangle=0\ ,$ (16)
since $\phi^{++}_{\text{np},p}\in\mathcal{L}_{++}$ and $\mathcal{L}_{++}$ is orthogonal to $\mathcal{L}_{--}$. The second-order perturbation theory correction is
@@
$=\langle\phi^{++}_{\text{np},p}|(-L_{--}V_{\text{i}}L_{--}+L_{++}V_{\text{i}}L_{--}-L_{--}V_{\text{i}}L_{++})\frac{1-L_{++}}{E^{++}_{\text{np},p}-H_{\text{np}}}$
$\quad\quad\quad\quad(-L_{--}V_{\text{i}}L_{--}+L_{++}V_{\text{i}}L_{--}-L_{--}V_{\text{i}}L_{++})|\phi^{++}_{\text{np},p}\rangle\$(18)
@@
where we used that $L_{--}H_{\text{np}}L_{--}=L_{--}(h_{1}+h_{2})L_{--}$.
@@
$H_{\text{wp}}=H_{\text{np}}+V_{\delta}=h_{1}+h_{2}+\left(L_{++}-L_{--}\right)V_{\text{i}}\ .$ (22)
By solving the with-pair equation, Eq. (S1), we obtain a resummation of the instantaneous $V_{\delta}$ double-pair effects. (It is necessary to note that only the double-pair corrections are summed up, but no crossed photon contributions are included in this way.)
$H_{\text{wp}}$ couples the $\mathcal{L}_{++}$ and $\mathcal{L}_{--}$ subspaces, but $H_{\text{wp}}$ is block diagonal over $\mathcal{L}_{++}\oplus\mathcal{L}_{--}$ and $\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$, which we exploit during the course of the numerical computations. Some mathematical properties regarding the block-structure of $H_{\text{wp}}$ are collected in the Supplementary Material. It is also shown there that although $H_{\text{wp}}$ in Eq. (S2) is non-hermitian, it can be rewritten to a hermitian form, and thus, it has real eigenvalues.
@@ III.1 Finite basis representation
where $\bm{p}=\bm{p}_{1}-\bm{p}_{2}$ is the relative momentum of the particles in the pBO case. The $X^{\mathrm{(p)BO}}$ balance was implemented as a metric ($X^{\mathrm{(p)BO}\dagger}X^{\mathrm{(p)BO}}$), and the Hamiltonian was transformed accordingly ($X^{\mathrm{(p)BO}\dagger}HX^{\mathrm{(p)BO}}$) Jeszenszki et al. (2021, 2022); Ferenc et al. (2022b, a); Mátyus et al. (2023).
@@
where $\left(I_{p}\right)_{q}=\delta_{pq}$. In the BO computations, we used explicitly correlated Gaussian (ECG) functions as spatial basis functions Szalewicz and Jeziorski (2010); Suzuki and Varga (1998); Mátyus and Reiher (2012); Mátyus (2019),
@@ III.2 Variational optimization of the basis functions
However, the basis sets that deliver us accurate no-pair (or non-relativistic) energies do not necessarily provide a reliable description of the $\mathcal{L}_{--}$ subspace required for the double-pair corrections. Moreover, since the with-pair Hamiltonian, Eq. (S2), is not bounded from below, the basis representation of the $\mathcal{L}_{--}$ subspace cannot be systematically improved through the minimization of the with-pair energy. In the next subsection, we first write the working formulae for the second-order perturbative corrections, and then, identify the quantity for which an auxiliary basis set is optimized to improve the $\mathcal{L}_{--}$ representation for the double-pair computations.
@@ III.3 Spectral representation of the perturbative corrections
The second-order perturbative correction, Eq. (20), for $V_{\text{i}}=V_{\rm C}+V_{\rm B}$ is written as
@@ IV.1 Comparison with nrQED through the $\alpha$ expansion
$\epsilon^{++,--}_{3,\text{CC}}(m_{1},m_{2})=-\frac{2}{3}\frac{\mu^{3}}{\pi}\left(\frac{2}{m_{1}^{3}}+\frac{1}{m_{1}m_{2}}+\frac{2}{m_{2}}\right)\;,$ (36)
@@
We note that this expression is analogous to Eq. (38), but instead of using the analytic result (with $m=m_{1}=m_{2}=1$) for $\langle\delta(\bm{r}_{12})\rangle_{\text{nr}}=\frac{m^{3}}{8\pi}$ ($n=1,l=0$, ground state), we compute $\langle\delta(\bm{r}_{12})\rangle_{\text{nr}}$ numerically.
@@ IV.1.2 Discussion of the numerical results
Tables 1 and 2 show the convergence of the second-order PT energies, Eqs. (30)–(33), for the double-pair instantaneous interaction corrections. The tables also show the $\varepsilon_{3}$ values obtained from repeating the computations for a series of $\alpha$ values and fitting the $\alpha$-dependent function of Eq. (34) to the data. These $\varepsilon_{3}$ values (free of higher-order $\alpha$ contributions) can be directly compared with the $\epsilon_{3}$ double-pair correction nrQED values, Eqs. (37)–(39) and (40)–(41).
@@ V Conclusions
$^{\text{c}}$ $\epsilon_{3}$ is the $\alpha^{3}E_{\mathrm{h}}$-order nrQED value, Eqs. (37)–eqrefeq:CBmm.
@@ Acknowledgements.
PJ gratefully acknowledges the support of the János Bolyai Research Scholarship of the Hungarian Academy of Sciences (BO/285/22). We thank the Momentum Programme of the Hungarian Academy of Sciences (LP2024-15/2024) and DKP for granting us access to the Hungarian Komondor HPC facility.
@@
Double-pair Coulomb(-Breit) photon correction to the correlated relativistic energy
@@
1 MTA–ELTE ‘Momentum’ Molecular Quantum electro-Dynamics Research Group, Institute of Chemistry, Eötvös Loránd University, Pázmány Péter sétány 1/A, Budapest, H-1117, Hungary
@@
$H_{\text{wp}}\Phi_{\text{wp},p}$ $=E_{\text{wp},p}\Phi_{\text{wp},p}\ ,$ (S1)
@@
First of all, let us consider the block structure of the with-pair Hamiltonian, Eq. (S2), over the two-particle $\mathcal{L}_{++}$ no-pair, the $\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$ Brown-Ravenhall (BR) Brown and Ravenhall (1951), and the $\mathcal{L}_{--}$ double-pair subspaces. For the interaction part, we have
@@
$L_{\text{BR}}$ labels the projector to the BR subspace. $\mathcal{L}_{\text{BR}}=\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$.
@@
$\left\langle\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] \phi^{\text{BR}}_{\mathrm{wp},i}\\[4.26773pt] -\phi^{--}_{\mathrm{wp},i}\end{array}\right.\Bigg|\quad\left.\Bigg|\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] 0\\[4.26773pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right\rangle\quad\text{for `electronic':}\quad|\phi^{++}_{\mathrm{wp},i}|^{2}>|\phi^{--}_{\mathrm{wp},i}|^{2}$ (S87)
@@
$\left\langle\begin{array}[]{@{}c@{}}-\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] \phi^{\text{BR}}_{\mathrm{wp},i}\\[4.26773pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right.\Bigg|\quad\left.\Bigg|\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] 0\\[4.26773pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right\rangle\quad\text{for `positronic':}\quad|\phi^{--}_{\mathrm{wp},i}|^{2}>|\phi^{++}_{\mathrm{wp},i}|^{2}$ (S94)

### v2

# Double-pair Coulomb and Breit photon correction to the correlated relativistic energy

@@ Abstract
The simplest, algebraic quantum-electrodynamical corrections due to the double-negative energy subspace and instantaneous interactions are computed to the no-pair energy of two-spin-1/2-fermion systems. Numerical results are reported for two-electron atoms with a clamped nucleus and positronium-like genuine two-particle systems. The Bethe-Salpeter equation provides the theoretical framework, and numerical methods have been developed for its equal-time time-slice. In practice, it requires solving a sixteen-component eigenvalue equation with a two-particle Dirac Hamiltonian, including the appropriate interaction. The double-pair corrections can either be included in the interaction part of the eigenvalue equation or treated as a perturbation to the no-pair Hamiltonian. The numerical results have an $\alpha$ fine-structure constant dependence that is in excellent agreement with the known $\alpha^{3}E_{\mathrm{h}}$-order double-pair correction of non-relativistic quantum electrodynamics.
@@ I Introduction
In a series of recent work Mátyus et al. (2023); Jeszenszki et al. (2021, 2022); Ferenc et al. (2022a); Ferenc and Mátyus (2023); Jeszenszki and Mátyus (2023); Margócsy and Mátyus (2024); Hollósy et al. (2024); Nonn et al. (2024), we have demonstrated that it should be possible, starting from the Bethe-Salpeter equation and its equal-time time-slice, to define a relativistic QED approach, in which a correlated relativistic reference state is first (numerically) computed to high precision Jeszenszki et al. (2021, 2022); Ferenc et al. (2022a); Ferenc and Mátyus (2023), and then, quantum electrodynamical corrections are computed to it by perturbation theory Mátyus et al. (2023); Margócsy and Mátyus (2024); Nonn et al. (2024). This second step is still in an early phase of research, and this work is one of the first steps to report actual methodologies and numerical values for the simplest QED corrections to a correlated relativistic reference state. This work considers the simplest, algebraic (energy-independent) correction. When further, important (but more complicated) corrections, i.e., due to retardation, multi-photon, and radiative, self-energy, vertex, and vacuum polarization ‘effects’ will be computed Mátyus et al. (2023); Margócsy and Mátyus (2024); Nonn et al. (2024), then, this direction will complement the already existing $1/Z$ and nrQED approaches, and may help solve the helium puzzle or provide independent tests for (parts of) the other approaches.
@@ II Double-pair ladder correction to the no-pair Dirac-Coulomb(-Breit) energy
$\Phi(\bm{r}_{1},\bm{r}_{2})$ is the equal-time time-slice of the two-particle wave function, depending only on the particles’ spatial coordinates, $\bm{r}_{1}$ and $\bm{r}_{2}$. $H_{\text{np}}$ is the no-pair Dirac-Coulomb(-Breit) (DC(B)) Hamiltonian and the $V_{\delta}+H_{\Delta}(E)$ term carries the corrections beyond the no-pair processes. The ‘no-pair’ term refers to excluding any interaction with the virtual electron–positron pairs, achieved by projecting the two-particle interaction onto the positive-energy subspace (vide infra). The present work focuses on instantaneous interactions; hence, we drop the energy-dependent $H_{\Delta}(E)$ term and will work with the algebraic $V_{\delta}$ contribution.
@@
$V_{\delta}$ $=(L_{++}-L_{--})V_{\text{i}}-L_{++}V_{\text{i}}L_{++}\;$
$=L_{++}V_{\text{i}}(1-L_{++})-L_{--}V_{\text{i}}\;$
$=L_{++}V_{\text{i}}L_{--}-L_{--}V_{\text{i}}L_{++}-L_{--}V_{\text{i}}L_{--}$
$+L_{++}V_{\text{i}}(L_{+-}+L_{-+})-L_{--}V_{\text{i}}(L_{+-}+L_{-+})\ .$ (7)
@@
where $i$ is the index of a general state. For solutions of the no-pair wave equation, Eq. (8), it is convenient to explicitly write out the contributions of $\Phi_{\text{np},i}$ according to the $\mathcal{L}_{++}$, $\mathcal{L}_{\text{BR}}=\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$, and $\mathcal{L}_{--}$ subspaces,
@@
in other words, $\phi_{\text{np},i}^{x}=L_{x}\Phi_{\text{np},i}L_{x}$ ($x=++,\text{BR}$ or --). The no-pair Hamiltonian is block diagonal over the $\mathcal{L}_{++},\mathcal{L}_{\text{BR}},$ and $\mathcal{L}_{--}$ subspaces. The non-trivial two-particle solutions, which correspond to the physically relevant bound and excited states, are in the $\mathcal{L}_{++}$ positive-energy subspace, hence we can focus on solving the eigenvalue equation on this subspace,
@@
where $E_{\text{np},m}^{--}=E_{\text{np},i}$ with $i\in\mathcal{I}_{--}$ and $\mathcal{I}_{--}$ collects the indexes of the negative-energy solutions of Eq. (8).
@@
$E_{p}^{(1)}$ $=\langle\phi^{++}_{\text{np},p}|V_{\delta}|\phi^{++}_{\text{np},p}\rangle=0\ ,$ (15)
where we used that $\phi^{++}_{\text{np},p}\in\mathcal{L}_{++}$ and $\mathcal{L}_{++}$ is orthogonal to $\mathcal{L}_{--}$ and the $\mathcal{L}_{\text{BR}}$ subspaces. The second-order perturbation theory correction is
@@
$=\langle\phi^{++}_{\text{np},p}|(L_{++}V_{\text{i}}L_{--}+L_{++}V_{\text{i}}L_{+-}+L_{++}V_{\text{i}}L_{-+})$
$\quad\quad\quad\quad\frac{1-L_{++}}{E^{++}_{\text{np},p}-H_{\text{np}}}(-L_{--}V_{\text{i}}L_{++})|\phi^{++}_{\text{np},p}\rangle$ (17)
@@
where we used the orthogonality of the $\mathcal{L}_{++},\mathcal{L}_{--},\mathcal{L}_{\text{BR}}$ subspaces, and in the last step, $L_{--}H_{\text{np}}L_{--}=L_{--}(h_{1}+h_{2})L_{--}$.
@@
$H_{\text{wp}}=H_{\text{np}}+V_{\delta}=h_{1}+h_{2}+\left(L_{++}-L_{--}\right)V_{\text{i}}\ ,$ (21)
where the contribution of the virtual pairs through the instantaneous interaction is also considered, hence the short name, ‘with-pair’. By solving the with-pair equation, Eq. (S1), we obtain a resummation of the instantaneous $V_{\delta}$ double-pair effects. (It is necessary to note that only the double-pair corrections are summed up, but no crossed-photon contributions are included in this way.)
It is shown in the Supplementary Material that although $H_{\text{wp}}$ couples the $\mathcal{L}_{++}$ and $\mathcal{L}_{--}$ subspaces, it is block diagonal over $\mathcal{L}_{++}\oplus\mathcal{L}_{--}$ and $\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$, which we can exploit during the course of the numerical computations. Some further mathematical properties regarding the block-structure of $H_{\text{wp}}$ are collected in the Supplementary Material. It is also shown there that although $H_{\text{wp}}$ in Eq. (S2) is non-hermitian, it can be rewritten to a hermitian form, and thus, it has real eigenvalues.
@@ III.1 Finite basis representation
where $\bm{p}=\bm{p}_{1}-\bm{p}_{2}$ is the relative momentum of the particles in the pBO case and the superscript [4] indicates the matrix dimension. The 1 and 2 subscripts of $\bm{\sigma}^{[4]}$ are the particle indices. They are defined as $\bm{\sigma}_{1}^{[4]}=\bm{\sigma}^{[2]}\otimes 1^{[2]}$ and $\bm{\sigma}_{2}^{[4]}=1^{[2]}\otimes\bm{\sigma}^{[2]}$, originating from the block-wise direct product of the single-particle Dirac Hamiltonians in the two-particle theory. The $X^{\mathrm{(p)BO}}$ balance was implemented as a metric ($X^{\mathrm{(p)BO}\dagger}X^{\mathrm{(p)BO}}$), and the Hamiltonian was transformed accordingly ($X^{\mathrm{(p)BO}\dagger}HX^{\mathrm{(p)BO}}$) Jeszenszki et al. (2021, 2022); Ferenc et al. (2022b, a); Mátyus et al. (2023).
@@
where $\left(\bm{I}_{q}\right)_{p}=\delta_{qp}$. In the BO computations, we used explicitly correlated Gaussian (ECG) functions as spatial basis functions Szalewicz and Jeziorski (2010); Suzuki and Varga (1998); Mátyus and Reiher (2012); Mátyus (2019),
@@ III.2 Variational optimization of the basis functions
However, the basis sets that deliver us accurate no-pair (or non-relativistic) energies do not necessarily provide a reliable description of the $\mathcal{L}_{--}$ subspace required for the double-pair corrections. Moreover, since the with-pair Hamiltonian, Eq. (S2), is not bounded from below, the basis representation of the $\mathcal{L}_{--}$ subspace cannot be systematically improved through the minimization of the with-pair energy. In the next subsection, we first write the working formulae for the second-order perturbation-theory (PT) corrections, and then, identify the quantity for which an auxiliary basis set is optimized to improve the $\mathcal{L}_{--}$ representation for the double-pair computations.
@@ III.3 Spectral representation of the perturbative corrections
The second-order PT correction, Eq. (19), for $V_{\text{i}}=V_{\rm C}+V_{\rm B}$ is written as
@@ IV.1 Comparison with nrQED through the $\alpha$ expansion
$\epsilon^{++,--}_{3,\text{CC}}(m_{1},m_{2})=-\frac{2}{3}\frac{\mu^{3}}{\pi}\left(\frac{2}{m_{1}^{2}}+\frac{1}{m_{1}m_{2}}+\frac{2}{m_{2}^{2}}\right)\;,$ (35)
@@
Furthermore, the $\alpha^{4}\ln\alpha E_{\mathrm{h}}$ term for the Dirac-Coulomb case is also available from the literature Khriplovich et al. (1993); Ferenc and Mátyus (2023)
$\epsilon_{4,1,{\text{CC}}}^{--}(1,1)=-\frac{1}{16}\ .$ (38)
@@
We note that this expression is analogous to Eq. (38), but instead of using the analytic result (with $m=m_{1}=m_{2}=1$) for $\langle\delta(\bm{r}_{12})\rangle_{\text{nr}}=\frac{1}{8\pi}$ ($n=1,l=0$, ground state), we compute $\langle\delta(\bm{r}_{12})\rangle_{\text{nr}}$ numerically.
@@ IV.1.2 Discussion of the numerical results
Tables 1 and 2 show the convergence of the second-order perturbation-theory (PT) energies, Eqs. (29)–(32), for the double-pair instantaneous interaction corrections. The tables also show the $\varepsilon_{3}$ values obtained from repeating the computations for a series of $\alpha$ values and fitting the $\alpha$-dependent function of Eq. (33) to the data. These $\varepsilon_{3}$ values (free of higher-order $\alpha$ contributions) can be directly compared with the $\epsilon_{3}$ double-pair correction nrQED values, Eqs. (36)–(39) and (40)–(41).
@@ V Conclusions
$^{\text{c}}$ $\epsilon_{3}$ is the $\alpha^{3}E_{\mathrm{h}}$-order nrQED value, Eqs. (36)–(39).
@@
Supplementary Material
The Supplementary Material contains: (a) Mathematical properties of the eigenvalues and eigenfunctions of the with-pair Hamiltonian. (b) Calculation of $\epsilon_{3,{\rm CB}}^{--}$. (c) Convergence tables of the non-relativistic, no-pair Dirac-Coulomb, and with-pair Dirac-Coulomb energies. Further data are available through the Zenodo repository Jeszenszki and Mátyus (2025).
@@ Acknowledgements.
PJ gratefully acknowledges the support of the János Bolyai Research Scholarship of the Hungarian Academy of Sciences (BO/285/22). We thank the Momentum Programme of the Hungarian Academy of Sciences (LP2024-15/2024) and DKF for granting us access to the Hungarian Komondor HPC facility.
@@
Double-pair Coulomb and Breit photon correction to the correlated relativistic energy
@@
1 MTA–ELTE Lendület ‘Momentum’ Molecular Quantum electro-Dynamics Research Group, Institute of Chemistry, Eötvös Loránd University, Pázmány Péter sétány 1/A, Budapest, H-1117, Hungary
@@
$H_{\text{wp}}\Phi_{\text{wp},i}$ $=E_{\text{wp},i}\Phi_{\text{wp},i}\ ,$ (S1)
@@
First of all, let us consider the block structure of the with-pair Hamiltonian, Eq. (S2), over the two-particle $\mathcal{L}_{++}$ no-pair, the $\mathcal{L}_{\text{BR}}=\mathcal{L}_{+-}\oplus\mathcal{L}_{-+}$ Brown-Ravenhall (BR) Brown and Ravenhall (1951), and the $\mathcal{L}_{--}$ double-pair subspaces. For the interaction part, we have
@@
$L_{\text{BR}}$ labels the projector to the BR subspace.
@@
$\left\langle\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] \phi^{\text{BR}}_{\mathrm{wp},i}\\[1.42271pt] -\phi^{--}_{\mathrm{wp},i}\end{array}\right.\Bigg|\quad\left.\Bigg|\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] 0\\[1.42271pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right\rangle\quad\text{for `electronic':}\quad|\phi^{++}_{\mathrm{wp},i}|^{2}>|\phi^{--}_{\mathrm{wp},i}|^{2}$ (S87)
@@
$\left\langle\begin{array}[]{@{}c@{}}-\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] \phi^{\text{BR}}_{\mathrm{wp},i}\\[1.42271pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right.\Bigg|\quad\left.\Bigg|\begin{array}[]{@{}c@{}}\phi^{++}_{\mathrm{wp},i}\\[1.42271pt] 0\\[1.42271pt] \phi^{--}_{\mathrm{wp},i}\end{array}\right\rangle\quad\text{for `positronic':}\quad|\phi^{--}_{\mathrm{wp},i}|^{2}>|\phi^{++}_{\mathrm{wp},i}|^{2}$ (S94)